OpenStudy (anonymous):

2x-3[5x-2(6-x)]Use the distributive property to simplify the expression.

OpenStudy (mathstudent55):

Distributive property: \(a(b + c) = ab + ac\) \(a(b - c) = ab - ac\)

OpenStudy (mathstudent55):

Example: \(-3(x - 4) = (-3)(x) - (-3)(4) = -3x + 12\)